one banana and one pear costs 95 cent. two banana and one apple costs 120 cents. two pear and two apple costs 220 cent. what is the unit price of apple banana and pear ?
1
Expert's answer
2013-08-05T08:59:57-0400
banana= x; pear = y; apple = z

So:
x+y=95
2x+z=120
2y+2z=220| y+z=110

y=95-x
z=120-2x
95-x+120-2x=110

215-3x=110| 3x=105 | x=35
y=95-x| y=95-35=60
z=120-2x| z=120-2*35=50

Total:
banana35
pear60
apple50